Shutdown mrc SNF 2009-02-04 10:26:51: Diffusion Pump not working

cbaxter at snf.stanford.edu cbaxter at snf.stanford.edu
Sat Feb 7 00:11:00 PST 2009


John, Eric and I installed newly cleaned dp pump. The foreline
pressure is 10mt after the dp heat up to temp, opened
up dp gate valve and flowed 20sccm of o2, pressure rise is
back to normal  at 1.2mt. Run 20 mins of Ar plasma w/out
problem.




More information about the mrc-pcs mailing list